Lindon Company is the exclusive distributor for an automotive product that sells for $30.00 per unit and has a CM ratio of 30%. The company’s fixed expenses are $162,000 per year. The company plans to sell 20,200 units this year.Required: 1. What are the variable expenses per unit? (Round your per unit answer to 2 decimal places.) 2. What is the brea
Answer #1

Answer:

1. Variable expense per unit $21.00
2. Break-even point in units 18,000
Break-even point in dollar sales $540,000
3. Unit sales needed to attain target profit 26,000
Dollar sales needed to attain target profit $780,000
4. New Break-even point in unit sales 13,500
New Break-even point in dollar sales $405,800
Dollar sales needed to attain target profit $585,000

Explanation:

1. Variable expense per unit

Contribution margin ratio =30%

$30×30% =$9

$30-$9 =$21

2.

Break-even point in units = $162,000/9 =18,000

Break-even point in dollar sales= $162,000/30% =$540,000

3.

Units needed to attain target profit

=($72,000+$162,000)/9 =26,000 units

Dollars needed to attain target profit

= 26,000×$30= $780,000

4.

Variable expense reduced by $3.00 per unit

Contribution margin per unit =$9+$3 =$12

New Break-even point in unit sales

= $162,000/$12 =13,500 units

New Break-even point in dollar sales

= 13,500×$30 =$405,800

Dollar sales needed to attain target profit

= $12/$30 =40% contribution margin ratio

$72,000= sales×40% -$162,000

Sales = ($162,000+$72,000)/40%

Sales = $234,000/40%

Sales =$585,000

____×____

All the best